# Contrastive Retrieval Heads Improve Attention-Based Re-Ranking

### v1

@@ Appendix A Prompt Structure
[document 1] Can’t Help Falling in Love ”Can’t Help Falling in Love” is a pop ballad
@@ B.1 Retrieval Head Location
We list below the exact location of the top $8$ NIAH retrieval heads, QR heads (without calibration) and CoRe heads for Llama-3.1 8B and Mistral 7B. We format ($L$-$H$) as layer $L$ head $H$, for example, (77-1919) means layer $7$ head $19$.
@@
We found that for Llama-3.1 8B, our top CoRe heads overlaps with QR heads in $3$ heads, but does not overlap with NIAH retrieval heads. For Mistral 7B, CoRe heads overlap in $2$ heads with both QR heads and NIAH retrieval heads. In both models, CoRe heads appear in earlier layers compared to other heads.
@@ B.2 Effect of Contrastive Temperature
As discussed in Section 4, the contrastive temperature $t$ controls the sharpness of the head distribution. Figure 7 demonstrates the effect of the temperature on the distribution of CoRe heads within Mistral 7B. Lower temperature heavily penalizes the non-CoRe heads which leads to larger gap in the $S_{CoRe}$. Nonetheless, all levels of temperatures result in similar distribution of the top $CoRe$ heads with difference in only $1$ or $2$ heads.
@@ Appendix C Additional Experiment Results
Figure 8: Distribution of $S_{CoRe}$ for all heads in Granite-3.2 8B with temperature $t=0.001$. The top CoRe heads are in middle layers or late layers.
@@
C.1 Results on Granite-3.2 8B Model
In this subsection, we provide experimental results on Granite-3.2 8B model.
Figure 9 demonstrates the distribution of CoRe heads within Granite-3.2 8B model, and the specific CoRe heads in Granite-3.2 8B are:
(1919-11), (1717-2020), (1919-1919), (3434-$28$), (1717-$25$), (1717-77), (1919-44), (1919-$31$).
Beside the CoRe heads concentrated in the middle layers, many heads in the late layers also exhibit high $S_{CoRe}$ which can potentially be considered CoRe heads. Nevertheless, most of the top $8$ CoRe heads still lie in the middle layers for Granite, with an exception for head (3434-$28$).
Figure 9 shows the nDCG@10 score on DBPedia dataset for Granite-3.2 8B with different number of retrieval heads. Similar to other models, we observe that the attention signal from fewer number of heads achieves better re-ranking accuracy than the noisy aggregation over all heads (ICR), and the nDCG@10 score peaks with a small number of CoRe heads. Interestingly, there is negligible accuracy degradation across different number of retrieval heads. This trend indicates that most attention heads within Granite-3.2 8B do not contribute much noise to the aggregated attention signals. Hence, the activation of more attention heads has a minimal influence on the final re-ranking accuracy. Figure 11 provides a closer look at the top $8$ CoRe heads with comparison to ICR and QR-R. Again, CoRe-R with more than $3$ heads consistently outperforms ICR and QR-R.
Figure 10: Average nDCG@10 on BEIR benchmark for Granite-3.2 8B using attention signal from different number of top retrieval heads.
@@ C.2 Results with Customized Prompts
As mentioned in Section 5.4, all attention-based re-rankers underperform the retriever baseline in some datasets that require more task-specific instruction prompts. For example, we show that a customized prompt for duplicated question dataset Quora, i.e. Please identify question that has the exact same meaning with the following query. Query: can improve the re-ranking performance of attention-based re-rankers.

### v2

@@ Appendix A Prompt Structure
[document 1] Can’t Help Falling in Love“Can’t Help Falling in Love” is a pop ballad
@@ B.1 Retrieval Head Location
Prior work Wu et al. (2025) show that only the top 101510-15 heads demonstrate strong retrieval signals. Our empirical results on the optimal number of heads (Section C.6) support this observation, as the re-ranking accuracy peaks within the top-$10$ heads and starts to decline after $10$ heads. This indicates the top few heads are the most important heads for retrieval tasks, while heads outside of top-$10$ are less effective which may contribute noise to the retrieval process.
We list below the exact location of the top $8$ NIAH retrieval heads, QR heads and CoRe heads for Llama-3.1 8B and Mistral 7B. We format ($L$-$H$) as layer $L$ head $H$, for example, (77-1919) means layer $7$ head $19$.
@@
We found that for Mistral 7B, CoRe heads overlap in $2$ heads with both QR heads and NIAH retrieval heads. For Llama-3.1 8B, our top CoRe heads overlap with QR heads in $3$ heads, but do not overlap with NIAH retrieval heads. Notably, the common 3 heads of CoRe and QR are in the top-3 and these are the heads with highest retrieval score. This explains why there is a close gap between QR-R and CoRe-R in the Llama results. Nevertheless, due to the difference in the remaining 5 heads, CoRe-R shows superior performance in many datasets that exhibit high number of hard negatives. In both models, CoRe heads appear in earlier layers compared to other heads, allowing for better accuracy-efficiency tradeoff with layer pruning.
Baselines Mistral Top-32 Mistral Top-64 Llama Top-32 Llama Top-64
NIAH heads 19 35 16 30
QR heads 25 48 27 53
Table 6: Number of overlapping heads between CoRe heads and other baselines.
We also investigate the overlapping in top-32 and top-64 heads between CoRe heads and the baselines in Table 6. In both top-32 and top-64 heads, we found that CoRe heads overlaps in approximately 50% with NIAH and 75% with QR. However, as mentioned above, these lower-ranking heads are not important for retrieval tasks, and their retrieval scores are very low compared to the first top-8 heads.
@@ B.2 Effect of Contrastive Temperature
As discussed in Section 4, the contrastive temperature $t$ controls the sharpness of the head distribution. Figure 8 demonstrates the effect of the temperature on the distribution of CoRe heads within Mistral 7B. Lower temperature heavily penalizes the non-CoRe heads which leads to larger gap in $S_{CoRe}$. Nonetheless, all levels of temperatures result in similar distribution of the top CoRe heads.
@@
We also report the exact CoRe heads detected in Mistral 7B and Llama-3.1 8B with varying temperature below.
Llama-3.1 8B CoRe heads:
  • $t=0.1$: (1313-$18$), (1313-11), (1414-$13$), (1313-2121), (1414-$31$), (1313-$13$), (88-$11$), (1414-2020).
  • $t=0.01$: (1313-$18$), (1414-$13$), (1313-2121), (1414-$31$), (1313-11), (1414-2020), (1313-$13$), (1313-33).
  • $t=0.001$: (1313-2121), (1313-$18$), (1414-$31$), (1414-$13$), (1414-2020), (1313-33), (1313-11), (1313-$13$).
Mistral 7B CoRe heads:
  • $t=0.1$: (1919-99), (1818-33), (1616-$12$), (1818-0), (1616-$22$), (1818-$22$), (1515-11), (1515-88).
  • $t=0.01$: (1919-99), (1616-$12$), (1515-11), (1616-$22$), (1515-2121), (1212-$11$), (1818-0), (1818-$22$).
  • $t=0.001$: (1515-2121), (1515-11), (1616-$12$), (1515-77), (99-$26$), (1212-$11$), (1212-77), (1818-0).
@@ Appendix C Additional Experiment Results
C.1 Results on Granite-3.2 8B Model
Figure 10 shows the nDCG@10 score on DBPedia dataset for Granite-3.2 8B with different number of retrieval heads. Similar to other models, we observe that the attention signal from fewer number of heads achieves better re-ranking accuracy than the noisy aggregation over all heads (ICR), and the nDCG@10 score peaks with a small number of CoRe heads. Interestingly, there is negligible accuracy degradation across different number of retrieval heads. This trend indicates that most attention heads within Granite-3.2 8B do not contribute much noise to the aggregated attention signals. Hence, the activation of more attention heads has a minimal influence on the final re-ranking accuracy.
@@ C.2 Results with Customized Prompts
All attention-based re-rankers underperform the retriever baseline in some datasets that require more task-specific instruction prompts. For example, we show that a customized prompt for duplicated question dataset Quora, i.e. Please identify question that has the exact same meaning with the following query can improve the performance.
@@
C.3 Results on Layer Pruning
We report the layer pruning experiment for the baselines NIAH-R and QR-R in Figure 11. As reported in Section B.1, NIAH and QR contain many mid-layer heads, NIAH-R and QR-R also show efficiency benefit with 30-40% layer pruning. However, both baselines start to degrade much earlier (after 40%) compared to CoRe-R which continues to retain near-identical performance up to 50%. This observation makes intuitive sense because all CoRe heads lie in early to mid layers while NIAH and QR have some late-layer heads (layer 20 and after). Overall, CoRe-R not only delivers the highest re-ranking accuracy, but also shows the best accuracy-efficiency tradeoff thanks to its structural head location.
(a) Mistral 7B
(b) Llama-3.1 8B
Figure 11: Average re-ranking accuracy on BEIR benchmark with different levels of layer pruning.
C.4 Results on Multi-hop Tasks
We conduct additional experiment on two multi-hop tasks MuSiQue (Trivedi et al., 2022) and CLIPPER (Pham et al., 2025). We follow the same setup in Chen et al. (2025) for MuSiQue dataset and adopt the same setting in Zhang et al. (2025) for CLIPPER dataset.
Table 8: Re-ranking performance on multi-hop datasets.
Dataset Retriever Mistral 7B Llama-3.1 8B Phi-4 Granite-3.2 8B
MuSiQue (R@2) 37.9 40.0 41.9 43.9 44.5 45.9 46.7 40.9 45.9 46.3 45.7 46.1 46.8
MuSiQue (R@5) 49.2 53.9 54.4 56.2 57.0 58.3 58.9 54.8 57.3 58.5 56.4 57.2 58.4
CLIPPER (R@2) 5.1 24.2 25.9 26.5 27.8 29.0 29.8 26.5 28.3 28.9 26.6 26.7 27.5
CLIPPER (R@5) 22.0 43.0 45.1 45.7 45.7 48.8 48.6 46.5 47.2 47.9 46.1 44.6 47.1
Table 8 reports Recall@2 and Recall@5 for both multi-hop datasets with four language models. Overall, CoRe-R stands out with the best re-ranking accuracy in both tasks across all models with only one exception of CLIPPER Recall@5 in Llama-3.1 8B.
C.5 Results on Different Detection Datasets
To study the sensitivity of CoRe-R to the detection dataset, we conduct the re-ranking experiments using different CoRe heads detected using a subset of the MSMARCO dataset. We found that the new detected CoRe heads are nearly identical to the CoRe heads detected using NQ, with $7$ out of $8$ overlapping heads. Specifically, the exact head location is listed below.
Llama-3.1 8B CoRe heads:
  • Detected via NQ: (1313-$18$), (1313-11), (1414-$13$), (1313-2121), (1414-$31$), (1313-$13$), (88-$11$), (1414-2020).
  • Detected via MSMARCO: (1313-$18$), (1313-11), (1414-$13$), (1313-2121), (1414-$31$), (1616-11), (88-$11$), (1414-2020).
Table 9: nDCG@10 of CoRe-R on BEIR benchmark with different hard negatives data.
Dataset Retriever Mistral 7B Llama-3.1 8B
Baseline NQ MSMARCO NQ MSMARCO
TREC-COVID 63.1 73.8 74.5 75.7 75.6
NFCorpus 33.7 35.2 35.6 36.7 36.5
DBPedia 36.0 37.5 38.1 39.4 39.4
SciFact 71.3 74.7 73.6 75.1 75.8
SciDocs 22.5 19.5 19.0 20.2 20.0
FiQA 36.9 41.4 41.3 44.1 43.9
NQ 51.6 58.1 58.3 63.2 63.3
FEVER 85.5 88.7 88.3 88.4 88.7
Climate-FEVER 30.3 23.3 22.7 22.7 22.7
HotpotQA 62.9 73.4 73.0 73.8 73.9
Touche 24.0 26.2 27.4 26.4 26.3
MSMARCO 30.7 31.7 32.3 34.9 35.1
Quora 86.7 85.2 83.8 82.3 82.0
ArguAna 56.4 52.7 52.2 55.0 54.7
CQADupstack 44.3 41.6 41.4 43.9 43.9
Average 49.1 50.9 50.8 52.1 52.1
Mistral 7B CoRe heads:
  • Detected via NQ: (1515-2121), (1515-11), (1616-$12$), (1515-77), (99-$26$), (1212-$11$), (1212-77), (1818-0).
  • Detected via MSMARCO: (1515-2121), (1515-11), (1616-$12$), (1515-77), (99-$26$), (1212-$11$), (1919-99), (1818-0).
We further repeat the re-ranking experiment on the BEIR benchmark using the CoRe heads detected via MSMARCO. The main results in Table 9 show similar and consistent performance, confirming that CoRe heads do not heavily depend on the detection data.
C.6 Results on Different Number of CoRe Heads
We repeat the “number of heads” experiment from Figure 2 on three additional datasets NQ, NFCorpus and FiQA using two models Mistral 7B and Llama-3.1 8B. Figure 12 shows the re-ranking results in each setting, and we observe the same consistent trend where the accuracy peaks within the top-10 CoRe heads and gradually declines as the number of heads grows. This stability across diverse datasets supports our claim that CoRe requires no dataset-specific tuning and that a small, fixed number of heads suffices in practice.
(a) Mistral 7B on NQ
(b) Mistral 7B on NFCorpus
(c) Mistral 7B on FiQA
(d) Llama-3.1 8B on NQ
(e) Llama-3.1 8B on NFCorpus
(f) Llama-3.1 8B on FiQA
Figure 12: nDCG@10 with CoRe-R across all number of CoRe heads.